OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 20 = x2 − x? A.(−4, 0) and (−5, 0) B.(−4, 0) and (5, 0) C.(4, 0) and (−5, 0) D. (4, 0) and (5, 0)

OpenStudy (anonymous):

if you want to know the procedure, then all you need to do is set y=0, then solve the resulting equation for x: \(\large 0+20=x^2-x \) otherwise, you can just check with each of those choices holds true in the equation...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x-intercept = y is 0 20=x^2-x x^2-x-20=0 (x-5)(x+4)=0 x=5 or x=-4 B

OpenStudy (anonymous):

thank you